Consider a series RC circuit for which R1 = R2 = 1
Answer #1

Equivalent resistance

1/Req = 1/R1 + 1/R2 +1/R3

1/Req = 1/1 +1/1 +1/2

Req =0.4 Mohms

equivalent capacitance

1/Ceq =1/C1 +1/C2 =1/5 + 1/4

Ceq =2.22 uF

TIme Constant

T =Req Ceq =0.4*2.22

T=0.8889 s

Maximum Charge

Qmax=CeqV =(2.22*10-6)*30 =6.666666667*10-5 C

In a RC circuit charge as a function of time is given by

Q=Qmax[1-e-t/T] =(6.66666667*10-5)[1-e-10/0.8889]

Q=6.66658*10-5 C